OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14 x = −32 and x = −4 x = −4 and x = 0 x = 0 and x = 28 No solution

OpenStudy (anonymous):

@juliannescriv

jimthompson5910 (jim_thompson5910):

I'll get you started |x + 2| + 16 = 14 |x + 2| + 16-16 = 14-16 |x + 2| + 0 = -2 |x + 2| = -2 and I'll let you take over from here

OpenStudy (anonymous):

Not sure how to solve it tbh...

OpenStudy (anonymous):

no solution

jimthompson5910 (jim_thompson5910):

what does absolute value represent?

OpenStudy (anonymous):

Can you help with a few more?

OpenStudy (anonymous):

yeah just tag me